# Boundary layers, transport and universal distribution in boundary driven active systems

### v2

@@ 3 Steady state distribution
    In the athermal $(D=0)$ case, $M=1$ irrespective of the dynamical parameters. Thus $J_{0}\propto(\Delta P-\overline{Q_{0}+Q_{1}})=P_{1}^{-}-P_{0}^{+}$, implying that the direction of the current is determined entirely by the boundary conditions. The presence of thermal noise drastically alters the picture. For any $D>0$, either of the activity parameters $v,\omega$ can be tuned to access all values of $M\in(0,1)$. Consequently, for a fixed boundary condition with the only constraint $P_{1}^{-}<P_{0}^{+}$ i.e. $10<\frac{\Delta P}{Q_{0}+Q_{1}}<$ 1, we can always tune the parameters across $M^{*}=\frac{\Delta P}{Q_{0}+Q_{1}}$ and a current reversal takes place.
  • VI.
    The current reversal, originating either from change in the boundary conditions or tuning of the dynamical parameters, leads to a sign flip in the slope of the bulk density profile. The bulk magnetisation being proportional to the current also changes sign. This suggests that the current governs the nature of the bulk profiles and not the boundary conditions.
  • VII.
    In the zero current condition, e.g. in a 1D box from where particles cannot escape, maintaining a nonzero net boundary magnetisation induces a global density difference, $\Delta P=M(Q_{0}+Q_{1})$. This phenomenon is reminiscent of the Seebeck effect with coefficient $M$ and the boundry magnetisations mimicking heat baths.
  • VIII.
@@
    For $D>0$ interesting nonmonotonicities are present in the current. A maximal current occurs in two distinct conditions in the large $L$ limit. Defining $D_{a}=\frac{v^{2}}{2\,\omega}$, the expression of the current in the thermodynamic limit becomes, $J_{0}\approx-\frac{D+D_{a}}{L}\,[\Delta P-\sqrt{\frac{D_{a}}{D+D_{a}}}\,(Q_{0}+Q_{1})]$ with $M\approx\sqrt{\frac{D_{a}}{D+D_{a}}}\,$. We can see that the competing currents due to the density gradient and boundary magnetisations are both enhanced but in a different manner as $D$ and $D_{a}$ are increased. The extremal current for fixed activity parameters and changing thermal noise occurs at $\frac{\partial J_{0}}{\partial D}=0$, which translates to $M=M_{p}\equiv 2\,\frac{\Delta P}{Q_{0}+Q_{1}}$ under the condition $0<\frac{\Delta P}{Q_{0}+Q_{1}}<\frac{1}{2}\,$. This condition also allows for the current reversal, implying that an optimal current can be achieved and the direction can be controlled by tuning $D$ only.
@@
    On the other hand, the extremal current condition for changing activity parameters is given by $\frac{\partial J_{0}}{\partial D_{a}}=0$, which implies, $M=M_{a}\equiv\frac{\Delta P}{Q_{0}+Q_{1}}-\sqrt{(\frac{\Delta P}{Q_{0}+Q_{1}})^{2}-1}\,$. This occurs only when $\frac{\Delta P}{Q_{0}+Q_{1}}>1$, which does not allow for a current reversal. Thus the conditions for optimal current clearly distinguishes the role of thermal diffusivity and the effective diffusivity due to activity. For $\frac{1}{2}<\frac{\Delta P}{Q_{0}+Q_{1}}<1$ the current is monotonous for all parameter values and only a current reversal occurs at $M^{*}$.
[…]
Figure 2: Left panel: Density (solid line) and magnetisation (dashed line) profiles in the steady state of the boundary driven RTP. Here $v=1.0,D=1.0,w=1.0,L=10.0$ and $P_{0}=P_{1}=0.1,Q_{0}=Q_{1}=0.1$. Right panel: Current reversal as the tumble rate $\omega$ is tuned. Here $v=1.0,D=1.0,L=10.0$ and $P_{0}=0.075,P_{1}=0.125,Q_{0}=Q_{1}=0.05$.
@@
Figure 3: Nonmonotonicity of the steady state current. Left panel: Diffusivity $D$ is tuned. Here $v=1.0,w=0.5,L=200.0$, $P_{0}=Q_{0}=0.002,P_{1}=Q_{1}=0.004$. A current reversal also occurs at higher values of $D$. Right panel: Tumble rate $\omega$ is tuned. Here $v=1.0,D=1.0,L=200.0$, $P_{0}=0.002,P_{1}=0.006,Q_{0}=Q_{1}=0.001$. (In both panels the $y$-axis is amplified $10^{4}$ times for better visibility.)
@@
It is surprising that this elementary model accommodates such remarkable out of equilibrium features, some of which are new according to our knowledge. These features might be general in the boundary driven active processes.
@@ 3.1 Steady state result for zero boundary magnetisation:
The equations for the steady state and the boundary conditions in this case are related to the exit probability $E_{\pm}(x)$ of the particle at the left boundary starting at an initial position $x$ with initial spin ±$\pm 1$. More details are in Appendix A. It gives a simulation scheme for the steady state. The data with zero boundary magnetisation is shown in Figure 4, which agrees with the theoretical expressions in Eqs. (10)-(11). We however couldn’t simulate the steady state with nonzero boundary magnetisations.
Figure 4: Steady state behaviour for $Q_{0}=Q_{1}=0~{}-$ (a): the probability density; (b): magnetisation for different tumble rate $\omega$. Here $L=10,v=1.0,D=1.0$ and the boundary densities are $P_{0}=0.15,P_{1}=0.02$. Points are obtained by simulating the exit probabilities and using Eq. (A.1).
@@ 4 Eigenspectrum and the time-dependent solution
$|\phi_{n}(x)\rangle=\alpha_{n}e^{k_{n}x}|A_{n}\rangle$ (20)
where $|A_{n}\rangle=[A_{n}^{+},A_{n}^{-}]^{T}$. Using the above, Eq. (19) can be written as,
@@
To facilitate the subsequent analysis we invert the above relation and express $k_{n}$ in terms of $\lambda_{n}$,
$k_{n}^{2}=\frac{\lambda_{n}}{D}+\frac{\omega\,D_{e}}{D^{2}}\pm\sqrt{\left(\frac{\lambda_{n}}{D}+\frac{\omega\,D_{e}}{D^{2}}\right)^{2}-\frac{\lambda_{n}^{2}+2\omega\lambda_{n}}{D^{2}}}~{},$ (24)
@@
$|A_{n}^{(i)}\rangle=\begin{bmatrix}a(k_{n}^{(i)})\\ 1\end{bmatrix}\,,i=1,...,4\,\,{\rm with}$ (26)
@@
Note that, $a(k_{n}^{(1)})\,a(k_{n}^{(2)})=a(k_{n}^{(3)})\,a(k_{n}^{(4)})=1$. Therefore, each eigenvalue $\lambda_{n}$ is four-fold degenerate and the corresponding eigenfunction $|\phi_{n}(x)\rangle$ can be written as a linear combination:
$|\phi_{n}(x)\rangle=\sum_{i=1}^{4}\alpha_{n}^{i}e^{k_{n}^{(i)}x}|A_{n}^{i}\rangle.$ (28)
@@
$|P(x,t)\rangle=|P_{ss}(x)\rangle+\sum_{n=1}^{\infty}c_{n}e^{\lambda_{n}t}\sum_{i=1}^{4}\alpha_{n}^{i}e^{k_{n}^{(i)}x}|A_{n}^{i}\rangle.$ (29)
The task is to determine $\lambda_{n}$’s and $\alpha_{n}^{i}$’s in terms of the parameters governing the dynamics, viz $v,\,\omega,\,D$ and the system size $L$. Recall the boundary conditions form Eq. (3),
@@
The time-dependent behaviour of the boundary driven case is therefore identical to that of the one-dimensional dynamics with two absorbing ends at $x=0$ and $L$. Henceforth, we shall omit the steady state. Using the absorbing boundary conditions in Eq.(28), we get the following set of equations for the coefficients $\alpha_{n}^{i}$ :
$x=0:~{}~{}$ $\sum_{i=1}^{4}\alpha_{n}^{i}=0~{},~{}~{}\sum_{i=1}^{4}\alpha_{n}^{i}a_{n}^{i}=0$ (32a)
$x=L:~{}~{}$ $\sum_{i=1}^{4}\alpha_{n}^{i}e^{k_{n}^{(i)}L}=0~{},~{}~{}\sum_{i=1}^{4}\alpha_{n}^{i}e^{k_{n}^{(i)}L}a_{n}^{i}=0$ (32b)
where $a_{n}^{i}\equiv a(k_{n}^{(i)})$. Above equations can be re-written as a matrix equation, $S(k_{n}^{(i)})|\alpha\rangle=0$, where $|\alpha\rangle=[\alpha_{n}^{(1)},\alpha_{n}^{(2)},\alpha_{n}^{(3)},\alpha_{n}^{(4)}]^{T}$ and
@@
To have a nontrivial solution for $\alpha$’s we must have
@@
which, upon solving, determines $\lambda_{n}$’s as a function of system parameters. We observe that, $\lambda=0,\,-2\omega$ satisfies the above determinant equation. $\lambda=0$ stands for the steady state which is just zero for the absorbing boundaries and is discarded. $\lambda=-2\omega$ corresponds to the tumble dynamics.
4.1 Symmetries of the model and solution for the spectrum
The determinant equation (34) is a complicated transcendental equation and we cannot find explicit solutions in general. However, the problem can be greatly simplified. Note that, the master equation (16) satisfies a symmetry, that is, whenever the spin is reversed ($\sigma\rightarrow-\sigma$) along with $x\rightarrow L-x$, the equation and the boundary conditions remain invariant. To express it mathematically, let us define an operation $\mathcal{O}_{x}$:
@@ 4.1 Symmetries of the model and solution for the spectrum
as shown in Figure 5. We find that, $\mathcal{O}_{x}$ commutes with $\mathcal{L}$, i.e. it is a symmetry operation. It also satisfies $\mathcal{O}_{x}^{2}=I$, implying that its eigenvalues are $o_{x}=\pm 1$. Consequently, the eigenstates $|\phi_{n}(x)\rangle$ of the operator $\mathcal{L}$ will also be an eigenstate of $\mathcal{O}_{x}$ corresponding to an eigenvalue ±$\pm 1$. In the following, we segregate all the eigenstates $|\phi_{n}(x)\rangle$ into two symmetry sectors labelled by the values of $o_{x}$, and determine the spectrum for each of the sectors.
@@
spectrum for $o_{x}=1$ (even sector):
Here, $\phi_{n}^{+}(x)=\phi_{n}^{-}(L-x)$. Using Eq. (28) and the values of $k_{n}$’s from Eq. (25), we find,
$\alpha^{(1)}\,a^{(1)}\,e^{k_{a}\,x}+\alpha^{(3)}\,a^{(3)}\,e^{k_{b}\,x}+\alpha^{(2)}\,a^{(2)}\,e^{-k_{a}\,x}+\alpha^{(4)}\,a^{(4)}\,e^{-k_{b}\,x}$ (36)
$=\alpha^{(1)}\,e^{k_{a}\,(L-x)}+\alpha^{(3)}\,e^{k_{b}\,(L-x)}+\alpha^{(2)}\,e^{-k_{a}\,(L-x)}+\alpha^{(4)}\,e^{-k_{b}\,(L-x)}.$ (37)
@@
$\alpha^{(1)}\,a^{(1)}=\alpha^{(2)}\,e^{-k_{a}\,L}$ (38)
$\alpha^{(3)}\,a^{(3)}=\alpha^{(4)}\,e^{-k_{b}\,L}$ (39)
Using the above expressions in Eqs. (32) for the coefficients $\alpha^{(i)}$, we obtain,
$\frac{\alpha^{(1)}}{\alpha^{(3)}}=-\frac{1+a^{(3)}e^{k_{b}\,L}}{1+a^{(1)}e^{k_{a}\,L}}=-\frac{a^{(3)}+e^{k_{b}\,L}}{a^{(1)}+e^{k_{a}\,L}}.$ (40)
@@
spectrum for $o_{x}=-1$ (odd sector):
Here, $\phi_{n}^{+}(x)=-\phi_{n}^{-}(L-x)$. Following the same procedure, we obtain the equation satisfied by the spectrum in this sector as,
@@
Although the equations (41)-(42) are exact and much simpler compared to Eq. (34), these are not yet exactly solvable for $\lambda$ in general. However, we earlier mentioned that $\lambda=0$ and $\lambda=-2\,\omega$ are eigenvalues of the evolution operator since these satisfy Eq. (34). Here we explicitly find that $\lambda=0$ satisfies Eq. (42) and therefor belongs to the odd sector (which, as argued earlier, should be discarded for these set of solutions). Whereas $\lambda=-2\,\omega$ satisfies Eq. (41) and belongs to the even sector. We find analytical expressions for the ‘band’ of eigenvalues close to 00 and $-2\,\omega$ in the large $L$ limit and the corresponding eigenfunctions in the two symmetry sectors. In the following, we demonstrate the results and discuss the relaxation behaviour.

### v4

@@ III. Current reversal and optimisation:
In the athermal $(D=0)$ case, $M$ equals $1$ irrespective of the dynamical parameters. Thus $J_{0}\propto(\Delta P-Q_{0}-Q_{1})=P_{1}^{-}-P_{0}^{+}$, implying that the direction of the current is determined entirely by the boundary conditions. The presence of thermal noise drastically alters the picture. For any $D>0$, $M=\frac{v}{\sqrt{2\,\omega D_{e}}}\leq 1$ in the large $L$ limit, and either of the activity parameters $v,\omega$ can be tuned to access all values of $M\in(0,1)$. Consequently, for a fixed boundary condition such that $P_{1}^{-}<P_{0}^{+}$ i.e. $0<\frac{\Delta P}{Q_{0}+Q_{1}}<1$, we can always tune $v,\omega$ in order to tune $M$ across $M^{*}=\frac{\Delta P}{Q_{0}+Q_{1}}$ leading to a reversal of the direction of the current. The two plots in Figure 2 correspond to different $\omega$ values such that the respective currents are oppositely directed, but with same boundary conditions. The same is possible by tuning $D$, as shown in the left panel of Figure 3.
@@
For $D>0$ interesting nonmonotonicities are present in the current. A maximal current occurs in two distinct conditions in the large $L$ limit. Defining $D_{a}=\frac{v^{2}}{2\,\omega}$, the expression of the current in the thermodynamic limit becomes, $J_{0}\approx-\frac{D+D_{a}}{L}\,[\Delta P-\sqrt{\frac{D_{a}}{D+D_{a}}}\,(Q_{0}+Q_{1})]$ with $M=\sqrt{\frac{D_{a}}{D+D_{a}}}\,$. We can see that the competing terms in the expression for $J_{0}$ are both enhanced but in a different manner as $D$ and $D_{a}$ are increased. The extremal current for fixed activity parameters and changing thermal noise occurs at $\frac{\partial J_{0}}{\partial D}=0$, which translates to $M=M_{p}\equiv 2\,\frac{\Delta P}{Q_{0}+Q_{1}}$ under the condition $0<\frac{\Delta P}{Q_{0}+Q_{1}}<\frac{1}{2}\,$. This condition also allows for the current reversal, implying that an optimal current can be achieved and the direction can be controlled by tuning $D$ only (Figure 3 left panel).
@@
On the other hand, the extremal current condition for changing activity parameters is given by $\frac{\partial J_{0}}{\partial D_{a}}=0$, which implies, $M=M_{a}\equiv\frac{\Delta P}{Q_{0}+Q_{1}}-\sqrt{(\frac{\Delta P}{Q_{0}+Q_{1}})^{2}-1}\,$. This occurs only when $\frac{\Delta P}{Q_{0}+Q_{1}}>1$, which does not allow for a current reversal (Figure 3 right panel). Thus the conditions for optimal current clearly distinguishes the role of thermal diffusivity and the effective diffusivity due to activity. For $\frac{1}{2}<\frac{\Delta P}{Q_{0}+Q_{1}}<1$ the current is monotonous for all parameter values and only a current reversal occurs at $M^{*}$.
It is useful to recall that the nature of the density profile depends crucially on the direction of the current – the latter is not determined solely by the boundary conditions, but can be controlled independently on tuning the dynamical parameters.
@@
Figure 3: Nonmonotonicity of the steady state current. Left panel: Diffusivity $D$ is tuned. Here $v=1.0,w=0.5,L=200.0$, $P_{0}=Q_{0}=0.002,P_{1}=Q_{1}=0.004$. A current reversal also occurs at higher values of $D$. Right panel: Tumble rate $\omega$ is tuned. Here $v=1.0,D=1.0,L=200.0$, $P_{0}=0.002,P_{1}=0.006,Q_{0}=Q_{1}=0.001$. (Note that in both panels the $y$-axis is amplified $10^{4}$ times for better visibility.)
@@
It is surprising that this simple model accommodates such remarkable out of equilibrium features, some of which are new according to our knowledge. These features might be general in the boundary driven active processes. The crucial role of the boundary magnetisation calls for a discussion of its possible realisations in experiments and simulations. This is nontrivial, in particular in the presence of boundary fluxes and translational noise, and is beyond the scope of this paper. We only outline a few possible avenues in the Appendix A.
@@ 3.1 Steady state result for zero boundary magnetisation:
The equations for the steady state and the boundary conditions in this case are related to the exit probability $E_{\pm}(x)$ of the particle at the left boundary starting at an initial position $x$ with initial spin $\pm 1$. More details are in Appendix B. It gives a simulation scheme for the steady state. We first simulate $E_{\pm}(x)$ for an RTP as follows: Consider an RTP with initial position $x\in[0,L]$ and initial spin $m=+1$; the dynamics stops as soon as the particle hits one of the boundaries. Counting the fraction of times it hits the left boundary gives $E_{+}(x)$. We similarly find $E_{-}(x)$ with m=1m=-1. The distribution is obtained thereafter using Eq. (B.1). The data with zero boundary magnetisation is shown in Figure 4, which agrees with the theoretical expressions in Eqs. (10)-(11). We however couldn’t simulate the steady state with nonzero boundary magnetisations.
Figure 4: Steady state for $Q_{0}=Q_{1}=0:$ (a) the probability density; (b) magnetisation for different tumble rate $\omega$. Here $L=10,v=1.0,D=1.0$ and the boundary densities are $P_{0}=0.15,P_{1}=0.02$. Points are obtained by simulating the exit probabilities and using Eq. (B.1).
@@ 4.1 Formalism
$|\phi_{n}(x)\rangle=\beta_{n}e^{k_{n}x}|A_{n}\rangle$ (20)
where $|A_{n}\rangle$ is a constant vector. Using the above, Eq. (19) can be written as,
@@
Inverting the above relation we express $k_{n}$ in terms of $\lambda_{n}$,
$k_{n}^{2}=\frac{\lambda_{n}}{D}+\frac{\omega\,D_{e}}{D^{2}}\pm\sqrt{\left(\frac{\lambda_{n}}{D}+\frac{\omega\,D_{e}}{D^{2}}\right)^{2}-\frac{\lambda_{n}^{2}+2\omega\lambda_{n}}{D^{2}}}~.$ (24)
@@
$|A_{n}^{(i)}\rangle=\begin{bmatrix}a(k_{n}^{(i)})\\ 1\end{bmatrix}\,,i=1,...,4\,\,~{\rm with}$ (26)
@@
Note that, $a(k_{n}^{(1)})\,a(k_{n}^{(2)})=a(k_{n}^{(3)})\,a(k_{n}^{(4)})=1$. Therefore, each eigenvalue $\lambda_{n}$ is four-fold degenerate and assuming that the $k_{n}^{(i)}$’s are all different, the eigenfunction $|\phi_{n}(x)\rangle$ can now be written as a linear combination:
$|\phi_{n}(x)\rangle=\sum_{i=1}^{4}\beta_{n}^{i}e^{k_{n}^{(i)}x}|A_{n}^{i}\rangle.$ (28)
@@
$|P(x,t)\rangle=|P_{ss}(x)\rangle+\sum_{n=1}^{\infty}c_{n}e^{\lambda_{n}t}\sum_{i=1}^{4}\beta_{n}^{i}e^{k_{n}^{(i)}x}|A_{n}^{i}\rangle.$ (29)
The task is to determine $\lambda_{n}$’s and $\beta_{n}^{i}$’s in terms of the parameters governing the dynamics, viz $v,\,\omega,\,D$ and the system size $L$. Recall the boundary conditions form Eq. (3),
@@
The time-dependent behaviour of the boundary driven case is therefore identical to that of the one-dimensional dynamics with two absorbing ends at $x=0$ and $L$. Henceforth, we shall omit the steady state. Using the absorbing boundary conditions in Eq.(28), we get the following set of equations for the coefficients $\beta_{n}^{i}$ :
$x=0:$~~ $\sum_{i=1}^{4}\beta_{n}^{i}=0~,~~\sum_{i=1}^{4}\beta_{n}^{i}a_{n}^{i}=0,$ (32a)
$x=L:$~~ $\sum_{i=1}^{4}\beta_{n}^{i}e^{k_{n}^{(i)}L}=0~,~~\sum_{i=1}^{4}\beta_{n}^{i}e^{k_{n}^{(i)}L}a_{n}^{i}=0,$ (32b)
where $a_{n}^{i}\equiv a(k_{n}^{(i)})$. Above equations can be re-written as a matrix equation, $S(k_{n}^{(i)})|\beta\rangle=0$, where $|\beta\rangle=[\beta_{n}^{(1)},\beta_{n}^{(2)},\beta_{n}^{(3)},\beta_{n}^{(4)}]^{T}$ and
@@
To have a nontrivial solution for $\beta$’s we must have
@@
which, upon solving, determines $\lambda_{n}$’s as a function of system parameters. We observe that, $\lambda=0,\,-2\omega$ satisfies the above determinant equation. $\lambda=0$ stands for the steady state which is just zero for the absorbing boundaries and is discarded. $\lambda=-2\omega$ is the relaxation rate of the tumble dynamics; it governs the relaxation of the initial magnetisation.
4.2 Symmetry of the model and solution for the spectrum
The determinant equation (34) is a complicated transcendental equation and we cannot find explicit solutions in general. However, the problem can be greatly simplified. Note that, the master equation (16) satisfies a reflection symmetry, that is, whenever the spin is reversed ($\sigma\rightarrow-\sigma$) along with $x\rightarrow L-x$, the equation and the boundary conditions remain invariant. To express it mathematically, let us define an operation $\mathcal{O}_{x}$
@@ 4.2 Symmetry of the model and solution for the spectrum
shown in Figure 5. We find that $\mathcal{O}_{x}$ commutes with $\mathcal{L}$, i.e. it is a symmetry operation.
@@
It satisfies $\mathcal{O}_{x}^{2}=I$, implying that its eigenvalues are $o_{x}=\pm 1$. Consequently, the eigenstates of $\mathcal{L}$ will also be an eigenstate of $\mathcal{O}_{x}$ corresponding to an eigenvalue $\pm 1$. In the following, we segregate all the eigenstates $|\phi_{n}(x)\rangle$ into two symmetry sectors labelled by the values of $o_{x}$, and determine the spectrum for each of the sectors.
In the $o_{x}=1$ (even) sector, $\phi_{n}^{+}(x)=\phi_{n}^{-}(L-x)$. Using Eq. (28) and the values of $k_{n}$’s from Eq. (25), we find,
$\beta^{(1)}\,a^{(1)}\,e^{k_{a}\,x}+\beta^{(3)}\,a^{(3)}\,e^{k_{b}\,x}+\beta^{(2)}\,a^{(2)}\,e^{-k_{a}\,x}+\beta^{(4)}\,a^{(4)}\,e^{-k_{b}\,x}$ (36)
$=\beta^{(1)}\,e^{k_{a}\,(L-x)}+\beta^{(3)}\,e^{k_{b}\,(L-x)}+\beta^{(2)}\,e^{-k_{a}\,(L-x)}+\beta^{(4)}\,e^{-k_{b}\,(L-x)}.$ (37)
@@
$\beta^{(1)}\,a^{(1)}=\beta^{(2)}\,e^{-k_{a}\,L}$ (38)
$\beta^{(3)}\,a^{(3)}=\beta^{(4)}\,e^{-k_{b}\,L}$ (39)
Using the above expressions in Eqs. (32) for the coefficients $\beta^{(i)}$, we obtain,
$\frac{\beta^{(1)}}{\beta^{(3)}}=-\frac{1+a^{(3)}e^{k_{b}\,L}}{1+a^{(1)}e^{k_{a}\,L}}=-\frac{a^{(3)}+e^{k_{b}\,L}}{a^{(1)}+e^{k_{a}\,L}}.$ (40)
@@
On the other hand, for the $o_{x}=-1$ (odd) sector, $\phi_{n}^{+}(x)=-\phi_{n}^{-}(L-x)$. Following the same procedure, we obtain the equation satisfied by the spectrum in this sector as,
@@
Although the equations (41)-(42) are exact and much simpler compared to Eq. (34), these are not yet exactly solvable for $\lambda$ in general. However, we earlier mentioned that $\lambda=0$ and $\lambda=-2\,\omega$ are eigenvalues of the evolution operator since these satisfy Eq. (34). Here we explicitly find that $\lambda=0$ satisfies Eq. (42) and therefor belongs to the odd sector (which, as argued earlier, should be discarded for these set of solutions). Whereas $\lambda=-2\,\omega$ satisfies Eq. (41) and belongs to the even sector. We find analytical expressions for the ‘band’ of eigenvalues close to 0 and $2\,\omega$ in the large $L$ limit and the corresponding eigenfunctions in the two symmetry sectors. In the following, we demonstrate the results and discuss the relaxation behaviour.